OpenStudy (anonymous):

What is the exact value of the expression sqrt 5 + 3 sqrt 10 + 2 sqrt 5: A. sqrt 20 B. 6 sqrt 20 C. 3 sqrt 5 + 3 sqrt 10 D. 2 sqrt 5 + 3 sqrt 10

OpenStudy (anonymous):

Combine like radicals. \[\sqrt{5} + 3\sqrt{10} + 2\sqrt{5} = 3\sqrt{5} + 3\sqrt{10}\]